OpenStudy (anonymous):

You have now written an equation where sin C equal to a fraction. Solve your equation for h.

OpenStudy (anonymous):

sin C = h/b

OpenStudy (danjs):

What is sin(C) = ?

OpenStudy (anonymous):

I don't know. The triangle doesn't show any values.

OpenStudy (danjs):

oh i see,, multiply both sides by b

OpenStudy (anonymous):

Explain please.

OpenStudy (danjs):

\[b * \sin(C) = \frac{ h }{ b }*b\]

OpenStudy (danjs):

now you have a b / b on the right, they cancel, left with just h